OpenStudy (anonymous):

Medal and fan for quick question please help! A firecracker shoots up from a hill 140 feet high, with an initial speed of 100 feet per second. Using the formula H(t) = -16t2 + vt + s, approximately how long will it take the firecracker to hit the ground? Five seconds Seven seconds Nine seconds 11 seconds

OpenStudy (anonymous):

well H(t) = -16t^2 + 100t + 140 then h(t) would equal to 0 0 = -16t^2 + 100t + 140 then solve for t

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@FriedRice .. Hey, Im having issues solving for t. i ran it threw www.mathway.com and gives me a weird answer =/

OpenStudy (anonymous):

hmmm ok

OpenStudy (anonymous):

god. sorry lol i cannot stand math.

OpenStudy (anonymous):

well you dont need mathway lol you can just plug in the points into the quadratic equation

OpenStudy (anonymous):

\[-b \pm \frac{ \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

-16t^2 + 100t + 140 a = -16 b = 100 c = 140

OpenStudy (anonymous):

\[-100 \pm \frac{ \sqrt{1000 -4(-16)(140)} }{ 2(-14) }\]

OpenStudy (anonymous):

so when i solve that it should then give me the answer?

OpenStudy (anonymous):

yeah but you might make a mistake so i'll just tell you the answer --------> 7.53357 or 7.5

OpenStudy (anonymous):

but you should try if you get that answer

OpenStudy (anonymous):

okay thank you so much @FriedRice !!!

OpenStudy (anonymous):

yw i wish i helped alot